OpenStudy (anonymous):

solve the polynomial equation: 15x^3-119x^2-10x+16=0

OpenStudy (campbell_st):

tough question, I'd graph it using some software or a calculator, to find the x-intercepts because just doing a quick check there is a root between 0 and 1, closer to 0

OpenStudy (anonymous):

x=8,x=-2/5,x=1/3

OpenStudy (campbell_st):

so how did you do it @pengembara_bumi1

OpenStudy (anonymous):

First thing you can do without a calculator is find all possible roots of the equation, which uses the leading coefficient (15) and the constant (16). Now we find all the factors of these: 15 has 1,3,5,15 16 has 1,2,4,8,16 Now place every leading coefficient factor over every constant factor and you have all of your possible choices. This one is quite a pain in the neck without a calculator because the only way to find it is to go through short division of every factor possible into the coefficients.

OpenStudy (anonymous):

just using calculator.. haha

OpenStudy (anonymous):

(x-8)(x+2/5)(x-1/3) Aren't the three constant terms supposed to multiply to give 16 though?

OpenStudy (campbell_st):

well remember that the coefficient of the leading term is 15 ... you cubic has a coefficient of the leading term being 1

OpenStudy (anonymous):

15 has 1,3,5,15 16 has 1,2,4,8,16 Looking at this, you have: ±1, 1/3, 1/5, 1/15, 2, 2/3, 2/5, 2/15, 4, 4/3, 4/5, 4/15, 8, 8/3, 8/5, 8/15, 16, 16/3, 16/5, and 16/15 as possible values.

OpenStudy (campbell_st):

so why not just graph it rather than substituting a silly number of possible values into an equation.

OpenStudy (anonymous):

I was always told to not use a calculator for some like these, including on tests that had upwards of 20 values to plug in without a calculator.

OpenStudy (campbell_st):

lol....my question to your teacher would be.... where is the higher learning in doing that

OpenStudy (campbell_st):

its interesting that you can use the rational root theorem but forgot about the actual equation... leading coefficient of 15.. but good luck with your maths